# A Fourier-Jacobi Dirichlet series attached to modular forms of $\textup{SO}(2,4)$

### v1

@@ 4. Dirichlet series
We will now focus on the $n=2$ case. Assume $k>3$ even and $(D,r)\in\widetilde{\textup{supp}}(L,\phi)$. We take $F\in S_{k}(\Gamma_{S})$ and consider the Dirichlet series
@@
$\mathcal{D}_{F,\mathcal{P}_{k,D,r}}(s):=\sum_{N=1}^{\infty}\langle\phi_{N},V_{N}P_{k,D,r}\rangle N^{-s},$ (11)
@@
where $\langle\ ,\ \rangle$ denotes the inner product on the space of Fourier–Jacobi cusp forms of weight $k$ and lattice index (see [Moc19, Definition 1.33]). As in [Psy25, Lemma 6.1], this converges absolutely and uniformly on compact subsets of $\mathbb{C}$ for $\textup{Re}(s)>k+1$.
@@
We can now rewrite $\mathcal{D}_{F,\mathcal{P}_{k,D,r}}(s)$ as in [Psy25, Proposition 7.3], but with a small difference. Let us make this explicit. Let
$\widehat{\Gamma}(L_{0}):=\{M\in\Gamma(L_{0})\mid M\in 1_{n+2}+\textup{M}_{n+2}(\mathbb{Z})S_{0}\},$ (12)
and write $\widehat{\Gamma}^{+}(L_{0})$ for the connected component of the identity inside $\widehat{\Gamma}(L_{0})$.
From now on, we choose $D=-\ell q$, with $\ell$ a positive integer and $q$ the level of $S$.
@@
We have that $\widehat{\Gamma}(L_{0})$ acts on $L_{0}\left[\ell,2^{-1}\mathbb{Z}\right]$ (see equation (4)) and there is a finite number of orbits under this action.
@@
We defer the proof of this to the next Section. ∎
Let now $h_{1}:=\#\left(L_{0}\left[\ell,2^{-1}\mathbb{Z}\right]/\widehat{\Gamma}(L_{0})\right)$. In a similar fashion to the discussion before [Psy25, Proposition 7.3], we fix a set of representatives $\{\xi_{i}\}_{i=1}^{h_{1}}$ for the orbits of this action, such that $\xi_{i}\in\mathcal{P}_{S}$ for all $i$. We also define
$n(\xi_{i};d):=\#\left\{s\in\mathbb{Z}^{2}/dS\mathbb{Z}^{2}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd},\,\,\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\right.\\ \left.\gamma\in\widehat{\Gamma}^{+}(L_{0})\right\}.$ (13)
We then have the following Proposition, as in [Psy25, Proposition 7.3].
@@
Let $(-\ell q,r)\in\widetilde{\textup{supp}}(L,\phi)$ and $\mathcal{P}$ be any finite set of primes, which includes the prime factors of $D$. We put
$\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}\langle V_{N}^{*}\phi_{N},P_{k,D,r}\rangle N^{-s},$
which converges absolutely for $\textup{Re}(s)>k+1$ by comparison with $D_{F,\mathcal{P}_{k,D,r}}(s)$. For each $i=1,\cdots,h_{1}$, let also
$\zeta_{\xi_{i},\mathcal{P}}(s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}n(\xi_{i};N)N^{-s}.$ (14)
We then have that
$\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)=\sum_{i=1}^{h_{1}}\zeta_{\xi_{i},\mathcal{P}}(s-k+3)D_{F,\xi_{i},\mathcal{P}}(s),$ (15)
where $D_{F,\xi_{i},\mathcal{P}}(s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\forall p\in\mathcal{P}\end{subarray}}^{\infty}A(N\xi_{i})N^{-s}$, and $h_{1}=\#\left(L_{0}\left[\ell,2^{-1}\mathbb{Z}\right]/\widehat{\Gamma}(L_{0})\right)$.
@@ 5. Equivalence of Quadratic Spaces
Our aim is now to understand the Dirichlet series $D_{F,\xi_{i},\mathcal{P}}(s)$. To that end, in this Section, we will introduce an isomorphism of the quadratic space $(V_{0},\phi_{0})$ with a quadratic space of Hermitian matrices over an imaginary quadratic field. Let
$S:=\begin{cases}\begin{pmatrix}2&0\\ 0&2m\end{pmatrix}\textup{ if }m\equiv 1,2\pmod{4}\\ \begin{pmatrix}2&1\\ 1&(m+1)/2\end{pmatrix}\textup{ if }m\equiv 3\pmod{4}\end{cases}$. (16)
for a square-free integer $m$, and set $K:=\mathbb{Q}\left(\sqrt{-\textup{det}S}\right)$. We note that when $K$ is of class number one (this is the case for which we will give our main result), by the theory of quadratic forms, there is only one class of quadratic forms with $\mathbb{Z}^{2}$ as a maximal lattice. That is, the choice of $S$ is canonical in the sense that any other choice would give us an isomorphic quadratic space, with the same arithmetic structure.
Let us now write $\mathcal{O}_{K}$ for the ring of integers of $K$. We have $\mathcal{O}_{K}=\mathbb{Z}[\omega]$, where
$\omega=\begin{cases}\sqrt{-m}\,\,\,\,\textup{ if }m\equiv 1,2\pmod{4}\\ \frac{1+\sqrt{-m}}{2}\,\,\,\,\textup{ if }m\equiv 3\pmod{4}\end{cases}$. (17)
We now consider the four dimensional (over $\mathbb{Q}$) quadratic space $(U,\psi)$, where
@@
and $\psi[h]=\det(h),\ \forall h\in U$. One has (see for example [Shi06b, (2.1d)]) that the corresponding bilinear form is given by $2\psi(x,y)=\textup{tr}(xy^{\iota})$, where $\iota$ here is the main involution on $M_{2}(K)$. We define the following elements of $U$
@@
and set $\Lambda:=\mathbb{Z}e+\mathbb{Z}f+\mathcal{O}_{K}g$. By [Shi06b, p. 495] we have that this is a $\mathbb{Z}$-maximal lattice in $U$.
@@
One sees easily that the map $f_{\omega}$ is well defined linear isomorphism of the $\mathbb{Q}$-vector spaces $U$ and $V_{0}$ (compare also with [Wer19, Lemma 1.24 and Lemma 1.25]). Moreover, an easy computation gives $\phi_{0}[f_{\omega}(u)]=\psi[u]$, for $u\in U$, and therefore $f_{\omega}$ is indeed an isometry. The last statement of the Lemma follows from Sylvester’s criterion. ∎
@@
$\displaystyle\begin{split}g_{\omega}:\textup{SO}^{\psi}(U)&\xrightarrow[]{\sim}\textup{SO}^{\phi_{0}}(V_{0})\\ \alpha&\longmapsto f_{\omega}\circ\alpha\circ f_{\omega}^{-1}\end{split}$ (19)
@@
with $a,b,c,d\in\mathbb{Q}$. From [Shi06b, (2.18)], we have that the condition on the ideals is equivalent to
$a\mathbb{Z}+d\mathbb{Z}+\textup{Tr}_{K/\mathbb{Q}}((b+\overline{\omega}c)\mathcal{O}_{K})=\mathbb{Z}.$ (20)
@@
$\phi_{0}(f_{\omega}(h),L_{0})=2^{-1}\mathbb{Z}\iff a\mathbb{Z}+d\mathbb{Z}-(2b+c)\mathbb{Z}-\left(b+\frac{m+1}{2}c\right)\mathbb{Z}=\mathbb{Z}.$ (21)
@@
$\phi_{0}(f_{\omega}(h),L_{0})=2^{-1}\mathbb{Z}\iff a\mathbb{Z}+d\mathbb{Z}-2b\mathbb{Z}-2mc\mathbb{Z}=\mathbb{Z}.$ (22)
@@
Again, it is now evident that (20) and (22) are equivalent. ∎
@@
$f_{\omega}(\gamma u\gamma^{*})=\Psi(\gamma)f_{\omega}(u),\ \forall\gamma\in\textup{SL}_{2}(\mathcal{O}_{K}),\ u\in U_{+},$ (23)
@@
From [Wer19, Theorem 3.17], we have that there is an isomorphism $\Phi:\textup{SU}(2,2;\mathcal{O}_{K})/\{\pm 1\}\longrightarrow\widehat{\Gamma}_{S}$. Moreover, there is an embedding $\epsilon:\textup{SL}_{2}(\mathcal{O}_{K})\xhookrightarrow{}\textup{SU}(2,2;\mathcal{O}_{K})$ by sending $\gamma\longrightarrow\textup{diag}(\gamma,\overline{\gamma}^{-t})$, which is compatible with the embedding $\eta:\widehat{\Gamma}^{+}(L_{0})\xhookrightarrow{}\widehat{\Gamma}_{S}$ via $\gamma\longmapsto\textup{diag}(1,\gamma,1)$, because of the explicit form of $\Phi$ given in [Wer19, Theorem 3.13]. We define a map $\Psi:\textup{SL}_{2}(\mathcal{O}_{K})/\{\pm 1\}\longrightarrow\widehat{\Gamma}^{+}(L_{0})$ by setting $\Psi(\gamma)$ the element such that $\Phi(\epsilon(\gamma))=\eta(\Psi(\gamma))$, and we claim this is an isomorphism with the required property. First of all, it is a homomorphism from the way it is defined. To check $\Psi$ is injective, assume $\Psi(\gamma_{1})=\Psi(\gamma_{2})$. Then $\Phi(\epsilon(\gamma_{1}))=\Phi(\epsilon(\gamma_{2}))$ and therefore $\epsilon(\gamma_{1})=\pm\epsilon(\gamma_{2})$, which in turn implies $\gamma_{1}=\pm\gamma_{2}$, as wanted. Now, if $M\in\widehat{\Gamma}^{+}(L_{0})$, we have $\eta(M)\in\widehat{\Gamma}_{S}$ and so there exists $\widetilde{\gamma}\in\textup{SU}(2,2;\mathcal{O}_{K})$ such that $\Phi(\widetilde{\gamma})=\eta(M)$. Now, let

### v2

@@ 6. Dirichlet Series
Assume now $\bm{F}\in S_{k}(D_{f})$ and $\bm{G}\in S_{k}(\widehat{D}_{f})$. For each fractional ideal $\mathfrak{m}$ of $\mathcal{O}_{K}$, consider the modular forms $F_{\mathfrak{m}},G_{\mathfrak{m}}$ defined by $F_{\mathfrak{m}}(Z):=\bm{F}(\widetilde{h}_{\mathfrak{m}};Z)$ and $G_{\mathfrak{m}}(Z):=\bm{G}(\widetilde{h}_{\mathfrak{m}};Z)$ for $Z\in\mathcal{H}_{S}$, with corresponding Fourier-Jacobi coefficients $\{\phi_{\mathfrak{m},N}\}_{N=1}^{\infty}$ and $\{\psi_{\mathfrak{m},N}\}_{N=1}^{\infty}$. For any finite set of primes $\mathcal{Q}$, we define the $\mathfrak{m}$-th Dirichlet series
@@
$D_{\mathcal{Q}}(F_{\mathfrak{m}},G_{\mathfrak{m}};s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\ \forall p\in\mathcal{Q}\end{subarray}}^{\infty}\langle\phi_{\mathfrak{m},N},\psi_{\mathfrak{m},N}\rangle N^{-s},$
@@
where $\langle\ ,\ \rangle$ is the inner product in the space of Fourier-Jacobi forms of weight $k$ and index $(L_{\mathfrak{m}},N\phi)$. By comparison with [psyroukis_orthogonal, Lemma 6.1], this Dirichlet series converges absolutely for $\textup{Re}(s)>k+1$ and represents a holomorphic function on this domain. We then set
(6.1) $D_{\mathcal{Q}}(\bm{F},\bm{G};s):=\sum_{\mathfrak{m}\in\textup{Cl}(K)}\zeta_{\mathfrak{m},\mathcal{Q}}(s-k+3)D_{\mathcal{Q}}(F_{\mathfrak{m}},G_{\mathfrak{m}};s),$
where
(6.2) $\zeta_{\mathfrak{m}}(s):=\sum_{\begin{subarray}{c}0\neq\mathfrak{a}\subset\mathcal{O}_{K}\\ \mathfrak{a}\in\mathfrak{m}\end{subarray}}N(\mathfrak{a})^{-s}$
denotes the zeta function attached to the ideal class $\mathfrak{m}$. Here $\mathfrak{m}$ runs over a set of representatives of the class group of $K$. This is also absolutely convergent and holomorphic for $\textup{Re}(s)>k+1$. We now consider the case $\bm{G}=\bm{P}$, the Poincaré series defined in Proposition 5.9. We then have that for each $\mathfrak{m}$ as above,
$D_{\mathcal{Q}}(F_{\mathfrak{m}},\mathcal{P}_{\mathfrak{m}};s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\ \forall p\in\mathcal{Q}\end{subarray}}^{\infty}\langle\phi_{\mathfrak{m},N},V_{N}P_{\mathfrak{m}}\rangle N^{-s}.$
Before we go any further, we make some remarks on how the Dirichlet series $D_{\mathcal{Q}}(F_{\mathfrak{m}},\mathcal{P}_{\mathfrak{m}};s)$ depends on the isometry class of the lattice $L_{\mathfrak{m}}$. In particular, we will show that it is independent of the choice of the representatives $\mathfrak{m}$.
Let us pick two lattices $L_{\mathfrak{m}_{1}}$ and $L_{\mathfrak{m}_{2}}$ and assume that there is exists an isometry $\sigma:L_{\mathfrak{m}_{1}}\longrightarrow L_{\mathfrak{m}_{2}}$. It is well known (see for example [jacobi_lattice, Section 4.1]) that if we extend $\sigma$ to obtain a bijection $\sigma:L_{\mathfrak{m}_{1}}\otimes_{\mathbb{Z}}\mathbb{C}\longrightarrow L_{\mathfrak{m}_{2}}\otimes_{\mathbb{Z}}\mathbb{C}$, then we have an isomorphism $S_{k}(L_{\mathfrak{m}_{2}},N\phi)\longrightarrow S_{k}(L_{\mathfrak{m}_{1}},N\phi)$, for any $N\geq 1$ given by $\psi(\tau,z)\longmapsto\sigma(\psi):=\psi(\tau,\sigma z)$. Moreover, we have that this map preserves the inner product, that is $\langle\psi_{1},\psi_{2}\rangle=\langle\sigma(\psi_{1}),\sigma(\psi_{2})\rangle$. This can be seen from the definition of the inner product as in [jacobi_lattice, Definition 1.33]. If we now write
$\psi(\tau,z)=\sum_{n\in\mathbb{Z},r\in L^{*}_{\mathfrak{m}_{2}}}c(n,r)e(n\tau+\phi(r,z)),$
then
$\sigma(\psi)(\tau,z)=\psi(\tau,\sigma(z))=\sum_{n\in\mathbb{Z},\ r\in L^{*}_{\mathfrak{m}_{1}}}c(n,\sigma(r))e(n\tau+\phi(r,z)).$
The isometry $\sigma$ induces a map $\sigma:L^{*}_{\mathfrak{m}_{1}}/L_{\mathfrak{m}_{1}}\longrightarrow L^{*}_{\mathfrak{m}_{2}}/L_{\mathfrak{m}_{2}}$, and hence a similar argument as in the proof of Lemma 5.8 gives that $P_{\mathfrak{m}_{1}}=\sigma(P_{\mathfrak{m}_{2}})$. Moreover, from the definition of the operator $V_{N}$, we see that $\sigma(V_{N}\psi)=V_{N}(\sigma(\psi))$ (see for example [jacobi_lattice, Theorem 4.26], for any Jacobi form $\phi$, and hence in particular $\sigma(V_{N}(P_{\mathfrak{m}_{2}}))=V_{N}(P_{\mathfrak{m}_{1}})$, for all $N\geq 1$.
Now we look at $\phi_{\mathfrak{m}_{i},N}$, for $i=1,2$, the $N$-th Fourier-Jacobi of the modular form $F_{\mathfrak{m}_{i}}$. We set $\sigma_{0}:=\textup{diag}(1,\sigma,1)\in G^{*}_{\mathbb{Q}}$. We then observe that if $\gamma:=h_{\mathfrak{m}_{1}}h^{-1}_{\mathfrak{m}_{2}}\sigma_{0}$, then $\gamma L_{\mathfrak{m}_{1},0}=L_{\mathfrak{m}_{1},0}$. In particular, $\gamma\in\Gamma(L_{\mathfrak{m}_{1},0})$, and so $\gamma_{1}:=\textup{diag}(1,\gamma,1)=\widetilde{h}_{\mathfrak{m}_{1}}\widetilde{h}_{\mathfrak{m}_{2}}^{-1}\textup{diag}(1,\sigma_{0},1)\in\Gamma(L_{\mathfrak{m}_{1},1})$. Hence, we have that $F_{\mathfrak{m}_{1}}|_{k}\gamma_{1}=F_{\mathfrak{m}_{1}}$, or equivalently $F_{\mathfrak{m}_{1}}(Z)=F_{\mathfrak{m}_{1}}(\gamma Z)=F_{\mathfrak{m}_{1}}(h_{\mathfrak{m}_{1}}h_{\mathfrak{m}_{2}}^{-1}\sigma_{0}(Z))$.
By Lemma 4.6, we have that
$F_{\mathfrak{m}_{1}}(Z)=\sum_{\eta\in L_{\mathfrak{m}_{1},0}^{*}}A(h^{-1}_{\mathfrak{m}_{1}}\eta)e(2\phi_{0}(\eta,h_{\mathfrak{m}_{1}}h_{\mathfrak{m}_{2}}^{-1}\sigma_{0}(Z)))=\sum_{\eta\in L_{\mathfrak{m}_{1},0}^{*}}A(h^{-1}_{\mathfrak{m}_{1}}\eta)e(2\phi_{0}(h_{\mathfrak{m}_{2}}h_{\mathfrak{m}_{1}}^{-1}\eta,\sigma_{0}(Z)))$
Setting $\eta\longmapsto h_{\mathfrak{m}_{1}}h^{-1}_{\mathfrak{m}_{2}}\eta^{\prime}$ with $\eta^{\prime}\in L^{*}_{\mathfrak{m},2}$ we obtain,
$F_{\mathfrak{m}_{1}}(Z)=F_{\mathfrak{m}_{2}}(\sigma_{0}(Z))$
But $\sigma_{0}(Z)=\sigma_{0}(\tau^{\prime},z,\tau)=(\tau^{\prime},\sigma(z),\tau)$, and hence we obtain that $\phi_{\mathfrak{m}_{1},N}(\tau,z)=\phi_{\mathfrak{m}_{2},N}(\tau,\sigma_{0}(z))$ or equivalently $\sigma(\phi_{\mathfrak{m}_{2},N})=\phi_{\mathfrak{m}_{1},N}$, for all $N\geq 1$.
That is, if $L_{\mathfrak{m}_{1}}$ is isometric to $L_{\mathfrak{m}_{2}}$, we then obtain
$D_{\mathcal{Q}}(F_{\mathfrak{m}_{1}},\mathcal{P}_{\mathfrak{m}_{1}};s)=D_{\mathcal{Q}}(F_{\mathfrak{m}_{2}},\mathcal{P}_{\mathfrak{m}_{2}};s).$
In particular, from the discussion in Section 4, we know that $L_{\mathfrak{m}_{1}}$ is isometric to $L_{\mathfrak{m}_{2}}$ if and only if the fraction ideal $\mathfrak{m}_{1}\mathfrak{m}_{2}^{-1}\in\textup{Cl}(K/\mathbb{Q})$. In particular, this is the case for when $\mathfrak{m}_{1}$ is in the same class as in $\mathfrak{m}_{2}$ in $\textup{Cl}(K)$. Since the zeta function $\zeta_{\mathfrak{m}}(s)$ is independent of the choice of representatives by definition, the Dirichlet series (6.1) is also independent of the choice of representatives. Moreover, we may write
$D_{\mathcal{Q}}(\bm{F},\bm{P};s)=\sum_{\mathfrak{m}\in\textup{Cl}(K)/\textup{Cl}(K/\mathbb{Q})}\left(\sum_{\mathfrak{n}\in\textup{Cl}(K/\mathbb{Q})}\zeta_{\mathfrak{m}\mathfrak{n},\mathcal{Q}}(s-k+3)\right)D_{\mathcal{Q}}(F_{\mathfrak{m}},P_{\mathfrak{m}};s).$
@@
We now rewrite $D_{\mathcal{Q}}(F_{\mathfrak{m}},P_{\mathfrak{m}};s)$ as in [psyroukis_orthogonal, Proposition 7.3], but with a small difference. Let us make this explicit.
First, for each $\mathfrak{m}$, it is clear that $\widehat{\Gamma}(L_{\mathfrak{m},0})$ acts on $L_{\mathfrak{m},0}\left[\ell,2^{-1}\mathbb{Z}\right]$ (see equation (2.3)) via left multiplication. By Lemma 4.4, we see that the map $v\longmapsto h_{\mathfrak{m}}v$, for $v\in L_{0}[\ell,2^{-1}\mathbb{Z}]$ induces a bijection $L_{0}\left[\ell,2^{-1}\mathbb{Z}\right]/\widehat{\Gamma}(L_{0})\longrightarrow L_{\mathfrak{m},0}\left[\ell,2^{-1}\mathbb{Z}\right]/\widehat{\Gamma}(L_{\mathfrak{m},0})$. We now have the following Lemma.
@@
For each $\mathfrak{m}$, there is a finite number of orbits of the action of $\widehat{\Gamma}(L_{\mathfrak{m},0})$ on $L_{\mathfrak{m},0}\left[\ell,2^{-1}\mathbb{Z}\right]$.
@@
From the above, it suffices to only show this for $\mathfrak{m}=\mathcal{O}_{K}$. We defer the proof of this just after Proposition 8.1. ∎
Hence, we may set $h_{1}:=\#\left(L_{\mathfrak{m},0}\left[\ell,2^{-1}\mathbb{Z}\right]/\widehat{\Gamma}(L_{\mathfrak{m},0})\right)$, independent of $\mathfrak{m}$. Moreover, in a similar fashion to the discussion before [psyroukis_orthogonal, Proposition 7.3], we fix a set of representatives $\{\xi_{i}\}_{i=1}^{h_{1}}$ for the representatives of $L_{0}\left[\ell,2^{-1}\mathbb{Z}\right]/\widehat{\Gamma}(L_{0})$, such that $\xi_{i}\in\mathcal{P}_{S}$ for all $i$. We then set $\xi_{\mathfrak{m},i}:=h_{\mathfrak{m}}\xi_{i}\in L_{\mathfrak{m},0}\left[\ell,2^{-1}\mathbb{Z}\right]$ for $i=1,\ldots h_{1}$. The set $\{\xi_{\mathfrak{m},i}\}$ is then a set of representatives for $L_{\mathfrak{m},0}\left[\ell,2^{-1}\mathbb{Z}\right]/\widehat{\Gamma}(L_{\mathfrak{m},0})$ and since we are taking $h_{\mathfrak{m}}\in G^{*,+}_{\mathbb{R}}$, we have that also $\xi_{\mathfrak{m},i}\in\mathcal{P}_{S}$.
We now define
(6.3) $n(\xi_{\mathfrak{m},i};d):=\#\left\{s\in\mathbb{Z}^{2}/dS_{\mathfrak{m}}\mathbb{Z}^{2}\mid D_{\mathfrak{m}}\equiv\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[s]\pmod{q_{\mathfrak{m}}d},\right.\\ \left.\left(\frac{\frac{1}{2}q_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}[s]-D_{\mathfrak{m}}}{q_{\mathfrak{m}}d},A_{\mathfrak{m}}S_{\mathfrak{m}}^{-1}s,d\right)^{t}=\gamma\xi_{\mathfrak{m},i},\gamma\in\widehat{\Gamma}^{+}(L_{\mathfrak{m},0})\right\}.$
We now set $\mathcal{P}:=\bigcup_{\mathfrak{m}}\mathcal{P}_{\mathfrak{m}}$, where $\mathcal{P}_{\mathfrak{m}}$ are finite sets containing all the prime factors of $D_{\mathfrak{m}}=-\ell q_{\mathfrak{m}}$. Using the expansions in (4.6), we then have the following Proposition, as in [psyroukis_orthogonal, Proposition 7.3].
@@
Let $(-\ell q_{\mathfrak{m}},r_{\mathfrak{m}})\in\widetilde{\textup{supp}}(L_{\mathfrak{m}},\phi)$. We then have
$D_{\mathcal{Q}}(F_{\mathfrak{m}},P_{\mathfrak{m}};s)=\sum_{i=1}^{h_{1}}\zeta_{\mathcal{Q}}(\xi_{\mathfrak{m},i};s-k+3)D_{\mathcal{Q}}(F_{\mathfrak{m}},\xi_{\mathfrak{m},i};s),$
where
$D_{\mathcal{Q}}(F_{\mathfrak{m}},\xi_{\mathfrak{m},i};s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\ \forall p\in\mathcal{Q}\end{subarray}}^{\infty}a(\widetilde{h}_{\mathfrak{m}};N\xi_{\mathfrak{m},i})N^{-s},$
and
(6.4) $\zeta_{\mathcal{Q}}(\xi_{\mathfrak{m},i};s):=\sum_{\begin{subarray}{c}d=1\\ (d,p)=1\ \forall p\in\mathcal{Q}\end{subarray}}^{\infty}n(\xi_{\mathfrak{m},i};d)d^{-s}.$
Using now Lemma 4.6, we can write $a(\widetilde{h}_{\mathfrak{m}};N\xi_{\mathfrak{m},i})=a(id,Nh_{\mathfrak{m}}^{-1}\xi_{\mathfrak{m},i})=a(id,N\xi_{i})=A(N\xi_{i})$ for all $N\not\in\mathcal{Q}$. In particular, we obtain that $D_{\mathcal{Q}}(F_{\mathfrak{m}},\xi_{\mathfrak{m},i};s)=D_{\mathcal{Q}}(F,\xi_{i};s)$ for all $\mathfrak{m}$.
We now select a set of representatives for $\textup{Cl}(K)$ to be relatively prime to any ramified ideals. That is, if we write $\mathfrak{d}_{K}$ for the different of $K$ over $\mathbb{Q}$, we assume that $(\mathfrak{m},\mathfrak{d}_{K})=1$. Therefore,
$D_{\mathcal{Q}}(\bm{F},\bm{P};s)=\sum_{\mathfrak{m}\in\textup{Cl}(K)}\zeta_{\mathfrak{m},\mathcal{Q}}(s-k+3)D_{\mathcal{Q}}(F_{\mathfrak{m}},P_{\mathfrak{m}};s)=\\ =\sum_{\mathfrak{m}\in\textup{Cl}(K)}\zeta_{\mathfrak{m},\mathcal{Q}}(s-k+3)\sum_{i=1}^{h_{1}}\zeta_{\mathcal{Q}}(\xi_{\mathfrak{m},i};s-k+3)D_{\mathcal{Q}}(F,\xi_{i};s)=\\ =\sum_{i=1}^{h_{1}}\left(\sum_{\mathfrak{m}\in\textup{Cl}(K)}\zeta_{\mathfrak{m},\mathcal{Q}}(s-k+3)\zeta_{\mathcal{Q}}(\xi_{\mathfrak{m},i};s-k+3)\right)D_{\mathcal{Q}}(F,\xi_{i};s),$
where
(6.5) $D_{\mathcal{Q}}(F,\xi_{i};s)=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\ \forall p\in\mathcal{Q}\end{subarray}}^{\infty}A(N\xi_{i})N^{-s}.$
@@ 7. Equivalence of Quadratic Spaces
Our aim is now to understand the Dirichlet series $D_{\mathcal{Q}}(F,\xi_{i};s)$. To that end, in this Section, we will introduce an isomorphism of the quadratic space $(V_{0},\phi_{0})$ with a quadratic space of Hermitian matrices over an imaginary quadratic field.
We consider the four-dimensional (over $\mathbb{Q}$) quadratic space $(U,\psi)$, where
@@
and $\psi[h]=\det(h),\ \forall h\in U$. One has (see for example [shimura2006quadratic, (2.1d)]) that the corresponding bilinear form is given by $2\psi(x,y)=\textup{tr}(xy^{\iota})$, where $\iota$ here is the main involution of $M_{2}(K)$. We define the following elements of $U$
@@
and set $\Lambda:=\mathbb{Z}e+\mathbb{Z}f+\mathcal{O}_{K}g$. By [shimura2006quadratic, p. 495] we have that this is a $\mathbb{Z}$-maximal lattice in $U$.
@@
One sees easily that the map $f_{\omega}$ is well defined linear isomorphism of the $\mathbb{Q}$-vector spaces $U$ and $V_{0}$ (compare also with [wernz, Lemma 1.24 and Lemma 1.25]). Moreover, an easy computation gives $\phi_{0}[f_{\omega}(u)]=\psi[u]$, for $u\in U$, and therefore $f_{\omega}$ is indeed an isometry. The last statement of the Lemma follows from Sylvester’s criterion. ∎
@@
(7.2) $\displaystyle\begin{split}g_{\omega}:\textup{SO}^{\psi}(U)&\xrightarrow[]{\sim}\textup{SO}^{\phi_{0}}(V_{0})\\ \alpha&\longmapsto f_{\omega}\circ\alpha\circ f_{\omega}^{-1}\end{split}$
@@
with $a,b,c,d\in\mathbb{Q}$. From [shimura2006quadratic, (2.18)], we have that the condition on the ideals is equivalent to
(7.3) $a\mathbb{Z}+d\mathbb{Z}+\textup{Tr}_{K/\mathbb{Q}}((b+\overline{\omega}c)\mathcal{O}_{K})=\mathbb{Z}.$
@@
(7.4) $\phi_{0}(f_{\omega}(h),L_{0})=2^{-1}\mathbb{Z}\iff a\mathbb{Z}+d\mathbb{Z}-(2b+c)\mathbb{Z}-\left(b+\frac{m+1}{2}c\right)\mathbb{Z}=\mathbb{Z}.$
@@
(7.5) $\phi_{0}(f_{\omega}(h),L_{0})=2^{-1}\mathbb{Z}\iff a\mathbb{Z}+d\mathbb{Z}-2b\mathbb{Z}-2mc\mathbb{Z}=\mathbb{Z}.$
@@
Again, it is now evident that (7.3) and (7.5) are equivalent. ∎
@@
(7.6) $f_{\omega}(\gamma u\gamma^{*})=\Psi(\gamma)f_{\omega}(u),\ \forall\gamma\in\textup{SL}_{2}(\mathcal{O}_{K}),\ u\in U_{+},$
@@
From [wernz, Theorem 3.17], we have that there is an isomorphism $\Phi:\textup{SU}(2,2;\mathcal{O}_{K})/\{\pm 1\}\longrightarrow\widehat{\Gamma}_{S}$. Moreover, there is an embedding $\epsilon:\textup{SL}_{2}(\mathcal{O}_{K})\xhookrightarrow{}\textup{SU}(2,2;\mathcal{O}_{K})$ by sending $\gamma\longrightarrow\textup{diag}(\gamma,\overline{\gamma}^{-t})$, which is compatible with the embedding $\eta:\widehat{\Gamma}^{+}(L_{0})\xhookrightarrow{}\widehat{\Gamma}_{S}$ via $\gamma\longmapsto\textup{diag}(1,\gamma,1)$, because of the explicit form of $\Phi$ given in [wernz, Theorem 3.13]. We define a map $\Psi:\textup{SL}_{2}(\mathcal{O}_{K})/\{\pm 1\}\longrightarrow\widehat{\Gamma}^{+}(L_{0})$ by setting $\Psi(\gamma)$ the element such that $\Phi(\epsilon(\gamma))=\eta(\Psi(\gamma))$, and we claim this is an isomorphism with the required property. First of all, it is a homomorphism from the way it is defined. To check $\Psi$ is injective, assume $\Psi(\gamma_{1})=\Psi(\gamma_{2})$. Then $\Phi(\epsilon(\gamma_{1}))=\Phi(\epsilon(\gamma_{2}))$ and therefore $\epsilon(\gamma_{1})=\pm\epsilon(\gamma_{2})$, which in turn implies $\gamma_{1}=\pm\gamma_{2}$, as wanted. Now, if $M\in\widehat{\Gamma}^{+}(L_{0})$, we have $\eta(M)\in\widehat{\Gamma}_{S}$ and so there exists $\widetilde{\gamma}\in\textup{SU}(2,2;\mathcal{O}_{K})$ such that $\Phi(\widetilde{\gamma})=\eta(M)$. Now, let